Multiply the following fractions.
2x^2 - x - 3 3x^2 - 11x - 20
___________ . _______________
3x^2+7x + 4 4x^2 - 9
Please help me thanks so much.

Answer by sdmmadam@yahoo.com(530)

(2x^2 - x - 3) (3x^2 - 11x - 20)
___________ . _______________
(3x^2+7x + 4) (4x^2 - 9)

=(2x-3)(x+1) (x-5)(3x+4)
___________ . _______________
(3x+4)(x+1) (2x+3)(2x-3)
=(x-5)/(2x+3) (canceling (2x-3)(x+1)(3x+4) )
Answer: (x-5)/(2x+3)
Note:(2x^2 - x - 3)
=2x^2-3x+2x-3
=x(2x-3)+1(2x-3)
=(2x-3)(x+1)
(3x^2 - 11x - 20)
=3x^2-15x+4x-20
=3x(x-5)+4(x-5)
=(x-5)(3x+4)
(3x^2+7x + 4)
=3x^2+4x+3x+4
=x(3x+4)+1(3x+4)
=(3x+4)(x+1)
(4x^2 - 9)
=(2x)^2-(3)^2
= (2x+3)(2x-3) (using a^2-b^2 = (a+b)(a-b)